OpenStudy (anonymous):

Rational Exponents I need to write this expression in the simplest form: (x^1/6 y^1/3)^-18 the -18 is loosing me please help

OpenStudy (anonymous):

Is it minus 18 or times -18

OpenStudy (anonymous):

its an exponent of -18

OpenStudy (anonymous):

so what does a negative exponent mean do you know?

OpenStudy (anonymous):

I know i divide but not sure how

OpenStudy (anonymous):

Negative exponents mean it is in the denominator

OpenStudy (anonymous):

So\[x^-2=1/x^2\]

OpenStudy (anonymous):

the answer should be 1 / x3 y6 just not sure how to get it

OpenStudy (anonymous):

What happens when you have a power raised to a power? You multiply the exponents right? So distribute the -18 and tell me what you get

OpenStudy (anonymous):

I think I just got it, I need to take 1/6*18/1 =3 and 1/3*18/1 =6 so i would get the answer just not sure if i did it the right way.

OpenStudy (anonymous):

Its not 18 its -18

OpenStudy (anonymous):

So you should get -3 and -6

OpenStudy (anonymous):

Now remember what i said earlier about negative exponents

OpenStudy (anonymous):

right but my class was just told if we have a negative rational exponent you take the negative away by putting it all over 1 so it would be 1 / (x1/6 y1/3) exponent 18 would then = 1 / x3 y6 if I did this right the answer matches the book answer

OpenStudy (anonymous):

yup that is it does it all make sense?

OpenStudy (anonymous):

I think so now that - was throwing me for a while but I realized it just needed to cross multiply to get my answer. I think I was making it harder than it is and talking it out helped.

OpenStudy (jhannybean):

Im not quite understanding what you're writing. \[(x^{1/6}y^{1/3})^{-18}\]Is thatyour question?

OpenStudy (anonymous):

yes that was the question

OpenStudy (jhannybean):

Oh, alright.

OpenStudy (anonymous):

In the future us (^) to denote raising something to a power

OpenStudy (anonymous):

*use

OpenStudy (anonymous):

how do i use that?

OpenStudy (anonymous):

Shift and 6

OpenStudy (anonymous):

ok thanks

OpenStudy (jhannybean):

So how I solved it: \[\large (x^{1/6}y^{1/3})^{-18} = \frac{1}{(x^{1/6}y^{1/3})^{18}} = \frac{1}{x^{(1/6) \cdot 18}y^{(1/3) \cdot 18}} =\frac{1}{x^3y^6} \]

OpenStudy (anonymous):

yes, thank you

OpenStudy (jhannybean):

Sorry, I was also curious as to what your question meant because i was cnofused a bit by the notation.